Master Fab Inc. is considering an investment in equipment that will replace direct labor. The equipment has a cost of $92,000 with a $8,000 residual value and a 10-year life. The equipment will replace one employee who has an average wage of $21,850 per year. In addition, the equipment will have operating and energy costs of $4,450 per year.

Determine the average rate of return on the equipment, giving effect to straight-line depreciation on the investment. If required, round to the nearest whole percent.
Answer #1

Average net income = 21850-4450-(92000-8000/10) = 9000

Average investment = 100000/2 = 50000

Average rate of return = average net income/Average investment = 9000/50000 = 18%